OpenStudy (anonymous):

HELP Please. Use implicit differentiation to find the equation of the tangent line to the curve xy^3+xy=4 at the point (2,1) . The equation of this tangent line can be written in the form: y=mx+b where M is: and where B is:

OpenStudy (dumbcow):

remember product rule for implicit \[(fg)' = f'g +fg'\]

OpenStudy (dumbcow):

M = slope = dy/dx

OpenStudy (anonymous):

oh so for implicit we use the product?

OpenStudy (anonymous):

always?

OpenStudy (dumbcow):

yes and chain rule....since it is with respect to x \[d/dx f(y) = f'(y)*\frac{dy}{dx}\]

OpenStudy (anonymous):

so which is f and which is g

OpenStudy (anonymous):

im a little confused since xy^3 are together.

OpenStudy (dumbcow):

depends on the term...first there is "xy^3" f(x) = x g(y) = y^3

OpenStudy (anonymous):

so f = x and g= y^3 ?

OpenStudy (anonymous):

so is the derivative: x*3y+ y^3*1 ?

OpenStudy (dumbcow):

yes, so result is \[1*y^{3}+ x*3y^{2} \frac{dy}{dx}\]

OpenStudy (anonymous):

dy/dx?

OpenStudy (dumbcow):

thats what the slope is...which is what we are trying to find

OpenStudy (anonymous):

all i know is the product rule is this: \[\frac{ d }{ dx } (fg)= f g \prime+g f \prime\]

OpenStudy (dumbcow):

anytime you differentiate a function of "y" wrt x you must multiply by "dy/dx" refer to my prev post

OpenStudy (anonymous):

otay. so then what.

OpenStudy (dumbcow):

repeat process for all terms of equation

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what do you mean

OpenStudy (dumbcow):

xy^3 +xy = 4 you have to differentiate each term, the "xy^3" and "xy" adn "4"

OpenStudy (anonymous):

3xy^2.....

OpenStudy (dumbcow):

?? first term is \[(y^{3} +3xy^{2} \frac{dy}{dx})\]

OpenStudy (dumbcow):

next term is "xy" .... derivative is \[(y + x \frac{dy}{dx})\] then of course , derivative of constant is 0 now you have \[y^{3} +3xy^{2} \frac{dy}{dx} +y+x \frac{dy}{dx} = 0\]

OpenStudy (dumbcow):

solve for "dy/dx" in terms of x and y

OpenStudy (anonymous):

okay i understand from here: \[\frac{ dy }{ dx}= 3xy ^{2}+ y ^{3}\]

OpenStudy (anonymous):

then from there, i don't know what to do.

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

well from there you would plug in the given point (2,1) to give you the slope (M) of tangent line but how did you solve for dy/dx .... your answer is wrong

OpenStudy (anonymous):

i didn't i'm doing this step by step

OpenStudy (anonymous):

from the result.. ^^ i just multiplied.

OpenStudy (dumbcow):

?? sorry i am not understanding

OpenStudy (anonymous):

remember before you wrote "yes, result is bla blah" then from there i just multiplied... x*(3y)^2 + y^3 *1

OpenStudy (anonymous):

you know.

OpenStudy (dan815):

hey mathcalculus

OpenStudy (dan815):

do you know chain rule

OpenStudy (dan815):